OpenStudy (anonymous):

If (2 − 3i) + (x + yi) = 6, what is x + yi?

OpenStudy (shubhamsrg):

Express the LHS in the form of A + iB, Now, if A + iB = p+ iq then, A=p and B =q Use this face to solve for x and y. Note that 6 can be written as 6 + 0i